Solve the differential equation of the following homogeneous equation. Show complete solution.


x dy/dx + y = x^3


1
Expert's answer
2021-10-03T17:16:32-0400

Let us solve the differential equation "x y'+ y = x^3," which is equivalent to the equation "(x y)' = x^3." It follows that "xy=\frac{x^4}4+C." Therefore, the general solution is of the form:

"y=\frac{x^3}4+\frac{C}x."